OpenStudy (anonymous):

compute the derivative y=(-6x^4+8x-2)^2

OpenStudy (anonymous):

chain rule the derivative of something squared is two times something, times the derivative of something

OpenStudy (anonymous):

o

OpenStudy (anonymous):

in math \[\frac{d}{dx}[f^2]=2ff'\]

OpenStudy (anonymous):

in your case \(f(x)=-6x^4+8x-2\), \(f'(x)=-24x^3+8\) and so your answer is \[2(-6x^4+8x-2)(-24x^3+8)\]